OpenStudy (anonymous):

Will give medal ------------------- Consider the following system of equations: y = −x + 2 y = 3x + 1 Which description best describes the solution to the system of equations? Line y = −x + 2 intersects line y = 3x + 1. Lines y = −x + 2 and y = 3x + 1 intersect the x-axis. Lines y = −x + 2 and y = 3x + 1 intersect the y-axis. Line y = −x + 2 intersects the origin.

OpenStudy (anonymous):

@mathwizzard3

OpenStudy (anonymous):

@ijlal

OpenStudy (ijlal):

the white line is of y=3x+1 the other is y=-x+2

OpenStudy (anonymous):

how do you make those graphs?

OpenStudy (ijlal):

they both do not intersect at origin you can see that in the graph so option 4 is wrong the rest 3 options are right but the best is option 1 that they both intersect each other so option 1 would be your answer for the best description of these equations :)

OpenStudy (ijlal):

Its a software called graphmatica download it from download.com

OpenStudy (anonymous):

ooh okay can you help me with another one please?

OpenStudy (anonymous):

Consider the following set of equations: Equation A: y = 2x + 4 Equation B: y = 3x + 1 Which of the following is a step that can be used to find the solution to the set of equations? 2x + 4 = 3x + 1 2x = 3x + 4 2x + 4 = 3x 2x + 1 = 3x + 4

OpenStudy (ijlal):

The first Step is correct 2x + 4 = 3x + 1 equate both equations to find the intersection point or in other words the solution set of the equation :)